Steinhoff Products, Inc., has a Sensor Division that manufactures and sells a number of products, including a standard sensor that could be used by another division in the company, the Safety Products Division, in one of its products. Data concerning that sensor appear below:

Steinhoff Products, Inc., has a Sensor Division that manufactures and sells a number of products, including a standard sensor that could be used by another division in the company, the Safety Products Division, in one of its products. Data concerning that sensor appear below:

Capacity in units 51,000
Selling price to outside customers $ 56
Variable cost per unit $ 37
Fixed cost per unit (based on capacity) $ 14

The Safety Products Division is currently purchasing 4,000 of these sensors per year from an overseas supplier at a cost of $48 per sensor.

Assume that the Valve Division is selling all of the valves it can produce to outside customers. From the standpoint of the Valve Division, what is the lost contribution margin if the valves are transferred internally rather than sold to outside customers?

Multiple Choice

  • $76,000

  • $969,000

  • $120,000

  • $20,000

The Safety Products Division is currently purchasing 4,000 of these sensors per year from an overseas supplier at a cost of $48 per sensor.

Assume that the Valve Division is selling all of the valves it can produce to outside customers. From the standpoint of the Valve Division, what is the lost contribution margin if the valves are transferred internally rather than sold to outside customers?

Multiple Choice

  • $76,000

  • $969,000

  • $120,000

  • $20,000
Answer #1

Lost Contribution margin is $20,000

Calculation :-

Contribution margin when selling to whole to outside custom :-

Sales - variable cost = contribution margin

(51,000 x 56) - (51,000 x 37) = $969,000

Contribution margin when valves are transferred internally :-

Internal transfer price = variable cost + recovery of fixed cost

= 37 + 14 = $51 per unit

Sales = (47,000 x 56) + (4,000 x 51) = 28,36,000

Contribution margin = 28,36,000 - (51,000 x 37) = $949,000

Contribution margin lost = 969,000 - 949,000 = $20,000